6. Four forces act on an object such that the object is at rest. Three of the forces are given by


F1 = 2i −2j, F2 = i −4j, F4 = −3i −5j. Determine F3 and its magnitude


1
Expert's answer
2022-05-13T11:15:44-0400

Let "F_3=ai+bj." Given the object is at rest


"F_1+F_2+F_3+F_4=0"

"x-axis: 2+1+a-3=0=>a=0"

"y-axis: -2-4+b-5=0=>b=11"

"F_3=11j."

"|F_3|=11"